what is the perimeter of a rectangle with an area of 33 2/9 ft squared and a length of 8 2/3 ft?
Answer by Cromlix(4381)

Hi there,
Convert 33 2/9 to ninths = 299/9
Convert 8 2/3 to thirds = 26/3
299/9 divided by 26/3
= 299/9 x 3/26
= 299/78 = 23/6
Width = 23/6 ft
Length = 26/3 ft
Perimeter = 2(Width) + 2(Length)
Perimeter = 2(23/6) + 2(26/3)
Perimeter = 23/3 + 52/3
Perimeter = 75/3
Perimeter = 25 ft.
Hope this helps :-)